OpenStudy (anonymous):

Maths Problems :) 1.)Mary and Ann together recieve a total of £30 for babysitting.Ann is paid for 14 days work and mary is payed for 15 days work.Ann's pay for 6 days work is £6 more than mary gets for 4 days. Work out how much they earn per day 2)Sam walks for x hours at 5km/h and runs for y hours at 10km/h.He travels a total of 35km and his average speed is 7km/h Find the value of x and y

OpenStudy (binary3i):

1) let one day payment of Ann be X and of Mary be Y so ' 14X+15Y=30 and other equation is 6X-6=4Y can you solve this?

OpenStudy (anonymous):

Nope not really im kinda new to this >_<

OpenStudy (binary3i):

total time he took is 5X+10Y average speed = total distance/ total time try it now. can you?

OpenStudy (anonymous):

yes, how would i finish the first question off?

OpenStudy (binary3i):

solve for X and Y

OpenStudy (binary3i):

you got it?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

the first question :(

OpenStudy (binary3i):

14X+15Y=30.........................................1 6X-6=4Y................................................2 from 2 y=3(x-1)/2 put it in 1 14x +15 *3(x-1)/2=30 28x+45x -45=60 73x=105 x=105/73=1.43 put it in 2 and get y y=3(x-1)/2 =3*.43/2=.65 x=1.43 y=.65

OpenStudy (anonymous):

according to my answers x=3 y=2 so how?? :(

OpenStudy (ghazi):

2) 5x+10y=35 and 7(x+y)=35 7x+7y =35 can you do it now?

OpenStudy (ghazi):

35x+50y=245 35x+35y=175 subtract equations 15y=70 y=70/15=14/3 and you can find x now :)

OpenStudy (binary3i):

total time is X+Y=35/7=5 total distance is 5x+10y=35 now when you solve this equations you get x=4,y=1

OpenStudy (ghazi):

hmm, cool :)

OpenStudy (anonymous):

how about question one then? ive got up to 14x+15y=30 and 6x-6=4y

OpenStudy (anonymous):

Ann should earn £9 and Mary should earn 12 :)

OpenStudy (ghazi):

correct

OpenStudy (anonymous):

how should i show my working ?

OpenStudy (ghazi):

14x+15y=30 6x-4y=6 -------------- 42x+45y=90 42x-28y=42 --------------- 73y=48 and find x i guess your solution and mine are different one

OpenStudy (anonymous):

Ghazi , after you finish this, please also help with my problem # thanks http://openstudy.com/study#/updates/50c791d7e4b0a14e4368569e

OpenStudy (binary3i):

no your wrong with 9 and 12. i have done above

OpenStudy (binary3i):

answers 1) x=1.43, y=.65 2) x=4,y=1

OpenStudy (anonymous):

nope, sorry thats wrong :)

OpenStudy (anonymous):

ghazi was correct x

OpenStudy (ghazi):

:D

OpenStudy (binary3i):

which one?

OpenStudy (anonymous):

gosh im confused and number 1

OpenStudy (anonymous):

here is my question http://openstudy.com/study#/updates/50c791d7e4b0a14e4368569e

OpenStudy (binary3i):

did you solved the equation

OpenStudy (anonymous):

nope